a stone of mass 20 g is released from a catapult whose rubber has been stretched through 4cm. if the force constant of the rubber is 200N/m calculate the velocity with which the phone leaves the catapult


1
Expert's answer
2021-05-14T18:22:51-0400

"kx^2/2=mv^2/2\to v=\sqrt{kx^2/m}=\sqrt{200\cdot 0.04^2/0.02}=4\ (m/s)" . Answer